Question 1: A) Kim just got a job. She is so happy that she decided to reward herself by purchasing a rare 1980 blue YUGO. TA-R-US used cars has given her the option of paying $14,000 now or making 8 annual payments of $2,000 starting in exactly 1 year. The interest rate is 9%. Which plan should Kim choose?


B) Emma’s goal in life is to own a ballet studio. Kenzie currently owns a studio, but she has plans to sell it. Emma agreed to pay Kenzie $3,000,000 5 years from today to acquire the studio. How much will Emma need to invest at 6% today to have enough money in 5 years?

C) When Charlie was paying for his gas at the gas station, he saw the shiny Kansas Lottery tickets and decided to play. To his surprise, he won $10,000! If Charlie invests his winnings in an account that earns him 6% interest each year, how much money will he have in 20 years?
Answer #1
A) Option 1 Pay now=$ 14000
Option 2 Making 8 annual payments of $2,000
Present value of option 2=2000*(PVA of $1 at 9% for 8 years)=2000*5.5348=$ 11070
Present value of option 2 is less than option 1
Hence,Kim should choose option 2
B) Amount need to invest=Present value=Amount required*(PV of $1 at 6% for 5 years)=3000000*0.7473=$ 2241900
C) Money will have in 20 years=Future value of investment=Investment*(FV of $1 at 6% for 20 years)=10000*3.2071=$ 32071